OpenStudy (anonymous):

If m=1-1/p and p-1=1/n^2, express m in terms of n only.

OpenStudy (ajprincess):

p-1=1/n^2 Solve for p. To do that u hav to add 1 both sides. Can u plz tell me vat u get?

OpenStudy (ajprincess):

@perederp

OpenStudy (anonymous):

@ajprincess , I've gotten p= 1+ 1/n^2

OpenStudy (ajprincess):

ya u r right:) Nw put 1+ 1/n^2 in place of p in m=1-1/p . Can u plz tell me vat u get @peredrep?

OpenStudy (anonymous):

@ajprincess , m = 1-1/ 1/ n^2

OpenStudy (ajprincess):

 Simplify it